OpenStudy (anonymous):

Use the discriminate to determine the number and type of solutions of the equation: -9 - 6x^2= -3x - 14

OpenStudy (anonymous):

this eqn simplifies to 6x^2-3x-5=0, discriminant is given by d=sqrt(b^2-4ac)(for a eqn ax^2+bx+c=0) now d=sqrt(-3)^2-4*6*(-5)=134

OpenStudy (anonymous):

if d=0,roots r real and equal d>0,roots r real and unequal d<o,roots r imaginary

OpenStudy (anonymous):

now since we have here a quadratic(degree is 2) eqn here.therefore no. of solutions of this eqn is 2 and since d>0 roots are real and unequal